# Small Quantum Low Parity Density Check Codes for Near-Term Experiments

### v1

@@ I Introduction
On the other hand, there has recently been a growing interest in quantum low-density parity-check (LDPC) codes which may have favorable scaling compared to the surface code and other local codes. It was mathematically shown that quantum LDPC codes that are asymptotically good exist [21, 22, 23, 24, 25, 26, 27, 28]. Here, a good code is a quantum code where the rate, meaning the number of logical qubits $k$ per physical qubits $n$, is constant as a function of the code distance $d$ as $d\rightarrow\infty$. However, real quantum devices do not operate in the asymptotic limit. Thus, the relevant question for near-term experiments is whether we can construct quantum LDPC codes with a smaller overhead than the surface code for small and intermediate scale quantum systems. An important step towards practical quantum LDPC codes was taken in Ref. [29], where a quantum code, known as a bivariate bicycle (BB) code, with the properties of $[[n{=}144,k{=}12,d{=}12]]$ was presented. We will refer to this code as the BB6-144, where the BB6 part signifies to the fact that the code features weight-6 stabilizers as opposed to the surface code which relies on weight-4 stabilizers. Similar codes were presented in Refs. [30, 31, 32, 33]. Additionally, in Ref. [34] it was shown that if the BB6-144 code is used in a fault tolerant architecture, it could reduce the total resources for practical algorithms by roughly one order of magnitude compared to a surface code architecture. On the other hand, there is still a number of practical challenges to realize a BB6-144 code with a real quantum device. Firstly, the code requires 288 physical qubits when including auxiliary qubits which is currently beyond most high-performing quantum devices that also support fast mid-circuit measurements. Secondly, any quantum LDPC code will naturally require long range interactions and it still remains an open question for many experimental platforms how to best implement these interactions. Another important aspect of any QEC code is that of decoding. The surface code can be efficiently decoded by the minimal-weight perfect-matching algorithm [35]. In contrast, a general quantum LDPC code requires a more involved decoding algorithm based, for example, on a belief propagation algorithm. Finally, weight-6 stabilizers of the BB6 codes imply that the connectivity for each qubit must be at least 6 which adds extra constraint on the device design. A first step towards real world experiments was presented in Ref. [36], where stabilizers were removed by hand such that the code becomes experimentally feasible. The error rate in this experiment was however not below the threshold for fault tolerance and the experiment did not yet demonstrate a clear advantage over surface codes.
@@ II Code construction
We begin by reviewing the framework of Ref. [30] which constructs trivariate bicycle (TB) codes. The key objects to construct these codes are the matrices $A$ and $B$ defined as a sum of $W_{A}$ and $W_{B}$ terms, respectively, where $W_{A}$ and $W_{B}$ are small integers, such that
@@
$k=2\,\text{dim}[\text{ns}(H_{X})\cap\text{ns}(H_{Z})],$ (8)
@@
where rs refers to the row-space and ||$\cdot$|| ⋅ | is the Hamming weight. An interesting observation is that the weight of the stabilizers in both $H_{X}$ and $H_{Z}$ are by construction equal to the integer number $W=W_{A}+W_{B}$, meaning the more terms we include in $A$ and $B$, the higher weight stabilizers we have.
@@
Here $Z_{i}$ and $X_{i}$ refers to the Pauli $Z$ and $X$ operators for data qubit $i$, respectively. Notice that there is a natural separation of the qubits into two halfs based on the construction of $H_{X}$ and $H_{Z}$. We refer to the first $n/2$ qubits as the left data qubits and the second half as the right data qubits. Each stabilizer connects, by construction, to exactly 2 left data qubits and 2 right data qubits. While this statement is true independent of the choice of $A$ and $B$ for fixed $W_{A}$ and $W_{B}$, we have picked here a code with the additional property that the left data qubits separate into two sets for the $Z$ stabilizers and the right data qubits separate into two sets for the $X$ stabilizers. This choice allows us to make a simple geometric toric layout of the code, see Fig. 1(a). In general for larger codes, this geometric construction is possible as long as we can separate each half into $l$ disconnected sets. At this point, we already see a modest improvement over surface codes as we only require 12 physical data qubits for 2 logical qubits. The original surface code of distance $d$ requires $d^{2}+(d-1)^{2}$ physical qubits per logical qubit meaning 25 data qubits are needed for a distance-3 code [47, 13]. On the other hand, the rotated version of the surface code requires only $d^{2}$ data qubits [48, 49]. Thus, when including the auxiliary qubits, the $[[12,2,3]]$ code presented here requires 24 physical qubits when including the auxiliary qubits while the rotated surface code requires 34 physical qubits for 2 distance-3 logical qubits. Other code constructions also exist with similar overhead, see for example Ref. [24]. As we will discuss in Sec. III.4, the reduced overhead compared to the surface code is not simply a constant scaling factor as for other local codes such as the color code [50, 51, 52, 12]. Rather, the code construction that we have reviewed here provides an improved scaling compared to local codes and via the code repository in Ref. [53] we provide the tool for anyone to construct a custom code using the TB code construction.
@@
As always, these logical operators are not unique as any multiplication of a stabilizers with the logical operator will yield an equivalent logical operator. For example, it is straight forward to notice that $X_{L1}=X_{4}X_{5}X_{6}$ is also a valid $X$-operator for the first logical qubit. Since the TB code that we construct here are of the Calderbank-Shor-Steane-type code, we can apply the logical $X$ and $Z$ operators transversally on the code. In Sec. IV, we will return to this point to explore how other logical gates can be applied within the codespace.
@@ III.1 Syndrome extraction circuit
Figure 4: Larger codes illustrated with the data qubits in blue, the $X$-type auxillarity qubits in orange and the $Z$-type auxiliary qubits in green. (a) A $[[24,4,3]]$ code constructed with $l=4$ and $m=3$. (b) A $[[56,4,5]]$ code constructed with $l=4$ and $m=7$. (c) A $[[88,4,7]]$ code constructed with $l=4$ and $m=11$.
@@ III.2 Decoding and error rates
As for the error extraction circuits, the simple weight-4 stabilizers also allow us to separate the $X$-type errors and the $Z$-type errors into two separate error graphs to decode similar to a surface code. As such, we can employ the minimal-weight perfect-matching algorithm (MWPM) [57, 58]. It is known that MWPM only approximates the error model since it essentially only considers X and Z errors [59]. However, with usage of several error correction rounds our code does not appear to be meaningfully hindered by this limitation. We use the implementation in Ref. [35] which takes the full error correction circuit implemented in stim [60] as input to construct the weights of the decoding problem. We implement the circuit as in Fig. 1(b) and we include a initialization error, measurement error and a depolarization error after all gates each with an error probability $P_{p}$ which we refer to as the physical error. A depolarization error on single qubit corresponds to the simultaneous application of $X$, $Y$ and $Z$ errors with equal strength. Following two-qubit gates, we depolarize the full two-qubit subspace. Using stim, we simulate the code for $d$ rounds of syndrome extraction before a final measurement of the data qubits and we decode the error syndromes to find the number of total errors $E$. Then, we can calculate the total logical error rate as $P_{k}=1-(1-E/N)^{1/r}$, where $N$ is the total number of repetitions and $r$ is the number of rounds. Finally, to readily compare codes with different number of qubits $k$, we extract the logical error rate per qubit $P_{L}=1-(1-P_{k})^{1/k}$, see Fig. 3(a) where we use $r=d$ for all simulations. In particular, we notice that at a physical error rate $0.2\times 10^{-3}$, the logical error per qubit is below $10^{-5}$. For comparison we also extract the logical error rate for a rotated surface code with the same error model and decoder, see Fig. 3(b). We notice a very similar performance as the $[[12,2,3]]$-code. In other words, the reduced overhead of the TB code does not lead to a decrease in the code performance.
@@ III.3 Larger codes
To demonstrate the performance of our code beyond a simple distance-3 code, we use the same construction as in Sec. II but increase $l$ and $m$. Specifically, we pick $l=4$ and $m=3$, $7$ and $11$ to construct codes with $k=4$ and distance $d=3$, $5$ and $7$, respectively. These codes each have their own characteristic choice of $A$ and $B$. For the distance 3 code with $k=4$ in Fig. 4(a), we have
@@
where $I$ is the identity matrix. We find that the performance per logical qubit matches fully with the $[[12,2,3]]$-code.
@@ III.4 Scaling
As shown in Fig. 5, the TB codes constructed here demonstrate a larger rate compared to the rotated surface code at the same code distance. More importantly, the scaling of $R$ with the distance $d$ is improved over the surface code. To quantify this, we performed a least-squares fit of the rate scaling with a power law of the form $R(d)=\alpha d^{-\beta}$ where $\alpha$ and $\beta$ are fit parameters. For the rotated surface code, $\beta\approx 2.0$ as expected, while for the TB codes we observe $\beta\approx 1.6$, showing a clear advantage from the additional non-local connections enabled by long-range couplers. This effectively allows us to encode more logical qubits per unit area without sacrificing distance to the same degree as a purely local construction.
It should be emphasized, however, that this result should not be over-interpreted, since there is no generic, systematic way of scaling the TB code to arbitrarily large distance $d$. For each target distance, the underlying matrices $A$ and $B$ in the code construction must be chosen by hand to satisfy the required distance and weight constraints. In particular, while in principle the quantum LDPC codes can yield large codes with finite rate and distance scaling sublinearly with $n$, practical realizations are limited by the feasibility of constructing and embedding the constituent codes in hardware, see also Sec. V. Nonetheless, the observed scaling exponent $\beta<2$ demonstrates the potential of architectures incorporating non-local couplers to overcome some of the rate limitations inherent to planar locality.
@@ V.2 Spin qubits based on electron shuttling
The use of electron spins in quantum dots as qubits was one of the early ideas for building quantum computers [88, 89]. In particular, semiconductor quantum devices are especially promising due to their scalable fabrication which is compatible with semiconductor foundries [90, 91, 92, 93, 94]. Moreover, recent developments have led to high-fidelity readout and gates [95, 96, 97, 98, 99, 100]. However, scaling has in practice been challenging due limited flexibility in the readout design as well as the lack of long-range interactions. To overcome these challenges, recent effort has been put into enabling coherent shuttling of electron (or hole) spin qubits across the device [40, 41, 42, 43, 44, 45, 46].
@@ VI Conclusion and Outlook
Following Ref. [30], we have introduced a family of trivariate bicycle (TB) codes, including a $[[12,2,3]]$ code and larger examples up to distance 7. These codes demonstrate favorable overhead compared to the rotated surface code while maintaining similar performance when decoded with a simple minimal-weight perfect-matching decoder. Our analysis reveals that the scaling of the code rate with distance improves over planar local codes. While the rotated surface code exhibits the expected $R\propto d^{-2}$ scaling, our TB constructions achieve a milder exponent of approximately $1.6$, illustrating how non-local connectivity can help alleviate the footprint constraints that limit strictly planar and local layouts. At the same time, we emphasize that these benefits come with the practical challenges of engineering long-range couplers or qubit shuttling schemes, and of constructing appropriate code parameters for each target distance.
@@
All the code and data to reproduce our results can be found at Ref. [53].
@@ Acknowledgments
The authors acknowledge the use of computational resources of the DelftBlue supercomputer, provided by Delft High Performance Computing Centre (https://www.tudelft.nl/dhpc). C.K.A. acknowledges support from the Dutch Research Council NWO.

### v3

# Small Quantum Low Density Parity Check Codes for Near-Term Experiments

@@ I Introduction
On the other hand, there has recently been a growing interest in quantum low-density parity-check (LDPC) codes which may have favorable scaling compared to the surface code and other local codes. It was mathematically shown that quantum LDPC codes that are asymptotically good exist [36, 68, 67, 17, 50, 24, 69, 51]. Here, a good code is a quantum code where the rate, meaning the number of logical qubits $k$ per physical qubits $n$, is always finite as a function of the code distance $d$ as $d\rightarrow\infty$. However, real quantum devices do not operate in the asymptotic limit. Thus, the relevant question for near-term experiments is whether we can construct quantum LDPC codes with a smaller overhead than the surface code for small and intermediate scale quantum systems. An important step towards practical quantum LDPC codes was taken in Ref. [15], where a quantum code, known as a bivariate bicycle (BB) code, with the properties of [[n=144,k=12,d=12]][[n{=}144,k{=}12,d{=}12]] was presented. We will refer to this code as the BB6-144, where the BB6 part signifies to the fact that the code features weight-6 stabilizers as opposed to the surface code which relies on weight-4 stabilizers. Similar codes with different requirements and constraints were presented in Refs. [94, 80, 101, 100]. Additionally, in Ref. [103] it was shown that if the BB6-144 code is used in a fault tolerant architecture, it could reduce the total resources for practical algorithms by roughly one order of magnitude compared to a surface code architecture. On the other hand, there is still a number of practical challenges to realize a BB6-144 code with a real quantum device. Firstly, the code requires 288 physical qubits when including auxiliary qubits which is currently beyond most high-performing quantum devices that also support fast mid-circuit measurements. Secondly, any quantum LDPC code will naturally require long range interactions and it still remains an open question for many experimental platforms how to best implement these interactions. Another important aspect of any QEC code is that of decoding. The surface code can be efficiently decoded by the minimal-weight perfect-matching algorithm [39]. In contrast, a general quantum LDPC code requires a more involved decoding algorithm based, for example, on a belief propagation algorithm. Finally, weight-6 stabilizers of the BB6 codes imply that the connectivity for each qubit must be at least 6 which adds extra constraint on the device design although modifications can be employed to reduce the required connectivity [80]. A first step towards real world experiments was presented in Ref. [95], where stabilizers were removed by hand such that the code becomes experimentally feasible. The error rate in this experiment was however not below the threshold for fault tolerance and the experiment did not yet demonstrate a clear advantage over surface codes.
@@ II Code construction
We begin by reviewing the framework of Ref. [94] which constructs general bicycle codes. The key objects to construct these codes are the matrices $A$ and $B$ defined as a sum of $W_{A}$ and $W_{B}$ terms, respectively, where $W_{A}$ and $W_{B}$ are small integers, such that
@@
$k=2\,\text{dim}[\text{ns}(A)\cap\text{ns}(B)],$ (8)
@@
where rs refers to the row-space and $|\cdot|$ is the Hamming weight. The code distance can be found through an efficient numerical optimization [80]. An interesting observation is that the weight of the stabilizers in both $H_{X}$ and $H_{Z}$ are by construction equal to the integer number $W=W_{A}+W_{B}$, meaning the more terms we include in $A$ and $B$, the higher weight stabilizers we have.
@@
Here $Z_{i}$ and $X_{i}$ refers to the Pauli $Z$ and $X$ operators for data qubit $i$, respectively. Notice that there is a natural separation of the qubits into two halfs based on the construction of $H_{X}$ and $H_{Z}$. We refer to the first $n/2$ qubits as the left data qubits and the second half as the right data qubits. Each stabilizer connects, by construction, to exactly 2 left data qubits and 2 right data qubits. While this statement is true independent of the choice of $A$ and $B$ for fixed $W_{A}$ and $W_{B}$, we have picked here a code with the additional property that the left data qubits separate into two sets for the $Z$ stabilizers and the right data qubits separate into two sets for the $X$ stabilizers. This choice allows us to make a simple geometric toric layout of the code, see Fig. 1(a). In general for larger codes, this geometric construction is possible as long as we can separate each half into $l$ disconnected sets. At this point, we already see a modest improvement over surface codes as we only require 12 physical data qubits for 2 logical qubits. The original surface code of distance $d$ requires $d^{2}+(d-1)^{2}$ physical qubits per logical qubit meaning 25 data qubits are needed for a distance-3 code [22, 26]. On the other hand, the rotated version of the surface code requires only $d^{2}$ data qubits [40, 90]. Thus, when including the auxiliary qubits, the $[[12,2,3]]$ code presented here requires 24 physical qubits when including the auxiliary qubits while the rotated surface code requires 34 physical qubits for 2 distance-3 logical qubits. This code is, incidentally, analogous to the twisted toric code of Ref. [17]. As we will discuss in Sec. III.4, the reduced overhead compared to the surface code is not simply a constant scaling factor as for other local codes such as the color code [13, 12, 31, 48]. Rather, the code construction that we have reviewed here provides an improved scaling compared to local codes and via the code repository in Ref. [5] we provide the tool for anyone to construct a custom code using the code construction from Ref. [94].
@@
As always, these logical operators are not unique as any multiplication of a stabilizers with the logical operator will yield an equivalent logical operator. For example, it is straight forward to notice that $X_{L1}=X_{4}X_{5}X_{6}$ is also a valid $X$-operator for the first logical qubit. Since the code that we construct here are of the Calderbank-Shor-Steane-type code, we can apply the logical $X$ and $Z$ operators transversally on the code. In Sec. IV, we will return to this point to explore how other logical gates can be applied within the codespace.
@@ III.1 Syndrome extraction circuit
Figure 4: Larger codes illustrated with the data qubits in blue, the $X$-type auxillarity qubits in orange and the $Z$-type auxiliary qubits in green. (a) A $[[24,4,3]]$ code constructed with $l=4$ and $m=3$. (b) A $[[56,4,5]]$ code constructed with $l=4$ and $m=7$. (c) A $[[104,4,7]]$ code constructed with $l=4$ and $m=13$.
@@ III.2 Decoding and error rates
As for the error extraction circuits, the simple weight-4 stabilizers also allow us to separate the $X$-type errors and the $Z$-type errors into two separate error graphs to decode similar to a surface code. As such, we can employ the minimal-weight perfect-matching algorithm (MWPM) [27, 28]. It is known that MWPM only approximates the error model since it essentially only considers X and Z errors [38]. However, with usage of several error correction rounds our code does not appear to be meaningfully hindered by this limitation. We use the implementation in Ref. [39] which takes the full error correction circuit implemented in stim [32] as input to construct the weights of the decoding problem. We implement the circuit as in Fig. 1(b) and we include a initialization error, measurement error and a depolarization error after all gates each with an error probability $P_{p}$ which we refer to as the physical error. A depolarization error on single qubit corresponds to the simultaneous application of $X$, $Y$ and $Z$ errors with equal strength. Following two-qubit gates, we depolarize the full two-qubit subspace. Using stim, we simulate the code for $d$ rounds of syndrome extraction before a final measurement of the data qubits and we decode the error syndromes to find the number of total errors $E$. Then, we can calculate the total logical error rate as $P_{k}=1-(1-E/N)^{1/r}$, where $N$ is the total number of repetitions and $r$ is the number of rounds. Finally, to readily compare codes with different number of qubits $k$, we extract the logical error rate per qubit $P_{L}=1-(1-P_{k})^{1/k}$, see Fig. 3(a) where we use $r=d$ for all simulations. In particular, we notice that at a physical error rate $0.2\times 10^{-3}$, the logical error per qubit is below $10^{-5}$. For comparison we also extract the logical error rate for a rotated surface code with the same error model and decoder, see Fig. 3(b). We notice a very similar performance as the $[[12,2,3]]$-code. In other words, the reduced overhead of the bicycle code does not lead to a decrease in the code performance.
@@ III.3 Larger codes
To demonstrate the performance of our code beyond a simple distance-3 code, we use the same construction as in Sec. II but increase $l$ and $m$. Specifically, we pick $l=4$ and $m=3$, 77 and $11$ to construct codes with $k=4$ and distance $d=3$, 55 and $7$, respectively. These codes each have their own characteristic choice of $A$ and $B$. For the distance 3 code with $k=4$ in Fig. 4(a), we have
@@
where $I$ is the identity matrix. We find that the performance per logical qubit matches fully with the $[[12,2,3]]$-code since this code is effectively a copy of two $[[12,2,3]]$-codes.
@@ III.4 Scaling
As shown in Fig. 5, the biclycle codes constructed here demonstrate a larger rate compared to the rotated surface code at the same code distance. More importantly, the scaling of $R$ with the distance $d$ is improved over the surface code. To quantify this, we performed a least-squares fit of the rate scaling with a power law of the form $R(d)=\alpha d^{-\beta}$ where $\alpha$ and $\beta$ are fit parameters. For the rotated surface code, $\beta\approx 2.0$ as expected, while for the quantum LDPC codes we observe $\beta\approx 1.7$, showing a clear advantage from the additional non-local connections enabled by long-range couplers. This effectively allows us to encode more logical qubits per unit area without sacrificing distance to the same degree as a purely local construction. Note, however, that the simple weight-4 connectivity used here means that all the constructed codes can be understood as surface code with twists and rotations. Thus, the reduced overhead comes directly from the from non-local connections required to form the appropriate modifications of the surface code [96].
It should be emphasized, however, that this result should not be over-interpreted, since there is no generic, systematic way of scaling a generalized bicycle code to arbitrarily large distance $d$. In particular, for weight-4 codes, any such generic scaling is bound to scale similar to the surface code [96]. For each target distance, the underlying matrices $A$ and $B$ in the code construction must be chosen by hand to satisfy the required distance and weight constraints. In particular, while in principle the quantum LDPC codes can yield large codes with finite rate and distance scaling sublinearly with $n$, practical realizations are limited by the feasibility of constructing and embedding the constituent codes in hardware, see also Sec. V. Nonetheless, the observed scaling exponent $\beta<2$ demonstrates the potential of architectures incorporating non-local couplers to overcome some of the rate limitations inherent to planar locality.
@@ V.2 Spin qubits based on electron shuttling
The use of electron spins in quantum dots as qubits was one of the early ideas for building quantum computers [53, 19]. In particular, semiconductor quantum devices are especially promising due to their scalable fabrication which is compatible with semiconductor foundries [107, 97, 62, 86, 30]. Moreover, recent developments have led to high-fidelity readout and gates [9, 89, 65, 99, 54, 61]. However, scaling has in practice been challenging due limited flexibility in the readout design as well as the lack of long-range interactions. To overcome these challenges, recent effort has been put into enabling coherent shuttling of electron (or hole) spin qubits across the device [29, 60, 104, 106, 92, 85, 41] which has also been proposed as a tool for surface codes [81] as well as for quantum LDPC codes [59].
@@ VI Conclusion and Outlook
Following Ref. [94], we have introduced a family of bicycle codes, including a $[[12,2,3]]$ code and larger examples up to distance 7. These codes demonstrate favorable overhead compared to the rotated surface code while maintaining similar performance when decoded with a simple minimal-weight perfect-matching decoder. Our analysis reveals that the scaling of the code rate with distance improves over planar local codes. While the rotated surface code exhibits the expected $R\propto d^{-2}$ scaling, our constructions achieve a smaller exponent of approximately $1.7$, illustrating how non-local connectivity can help alleviate the footprint constraints that limit strictly planar and local layouts. At the same time, we emphasize that these benefits come with the practical challenges of engineering long-range couplers or qubit shuttling schemes, and of constructing appropriate code parameters for each target distance.
@@
All the code and data to reproduce our results can be found at Ref. [5]. The distance finding algorithm has been adapted from Ref. [80].
@@ Acknowledgments
We acknowledge useful input from M. H. Shaw and B. M. Terhal. The authors acknowledge the use of computational resources of the DelftBlue supercomputer, provided by Delft High Performance Computing Centre (https://www.tudelft.nl/dhpc).
Funding Declaration
This research was supported by the European Union’s Horizon Europe programme under the Grant Agreement 101069515 (IGNITE). C.K.A. acknowledges support from the Dutch Research Council NWO.
Competing interests
The authors declare no competing interests.
Author Contributions
Both authors contributed to all aspects of conceptualizing, executing and writing of this work.